OpenStudy (anonymous):

HELP!!!!!!!!!! will fan and medal!!!!! Relationship B has a lesser rate than Relationship A. The graph represents Relationship A. Which equations could represent Relationship B? Choose all answers that are correct.

OpenStudy (anonymous):

We need the graph

OpenStudy (anonymous):

@sofiaalva

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@just_one_last_goodbye

OpenStudy (anonymous):

@YoungStudier

OpenStudy (just_one_last_goodbye):

omg Im so sorry xD

OpenStudy (just_one_last_goodbye):

I literally came and forgot about the question

OpenStudy (anonymous):

haha its fine

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

rate = slope so use the slope formula to find slope of the graph \[\huge\rm \frac{ y_2 -y_1 }{ x_2 -x_1 }\]

Nnesha (nnesha):

sub \[\large\rm (x_1,y_1)(x_2 ,y_2) ~~for ~~(1,3)(2,6)\]

OpenStudy (anonymous):

what?!?!?!?!?! i am so confused!!

Nnesha (nnesha):

rate means slope i gave u the slope formula plugin x 's and y's values

OpenStudy (anonymous):

ok?

Nnesha (nnesha):

(x,y) in this order pair first number represent x value and 2nd one y

OpenStudy (anonymous):

ok. so what do do?

Nnesha (nnesha):

then solve

Nnesha (nnesha):

here is an example if i have (3,4)(9,5) x_1 =3 x_2=9 y_1=4 y_2 =5 that's how i would plug in \[\frac{ 5-4 }{ 9-3}=\frac{1}{6}\]

Nnesha (nnesha):

plugin (1,3)(2,6) into the formula

OpenStudy (anonymous):

ok so B&C?

Nnesha (nnesha):

what is the slope of relationship A ??